If 10.3 g of magnesium nitride is treated with 10.3 g of water, what volume of ammonia gas would be collected at 24˚C and 0.989 atm?
1
Expert's answer
2020-06-15T14:44:14-0400

Mg3N2 + 6H2O --> 3Mg(OH)2 + 2NH3


n (Mg3N2) = 10.3g / 100.9g/mol = 0.102 mol

n (H2O) = 10.3g / 18.0g/mol = 0.572 mol

According to the equation, the Mg3N2 : H2O mole ratio is 1 : 6

The actual ratio is 1 : (0.572/0.102) = 1 : 5.6. Therefore H2O is the limiting reactant.

According to the equation, n(NH3) = n(H2O) / 3 = 0.191 mol


pV = nRT ==> V = nRT / p

T = 24oC = 297K

p = 0.989 atm = 100.2 kPa


V = (0.191mol x 8.31J/mol*K x 297K) / 100.2kPa = 4.7 L


Answer: 4.7 L